# A Novel Approach to Reduce Derivative Costs in Variational Quantum Algorithms

### v2

@@ Abstract
We present a detailed numerical study of an alternative approach, named Quantum Non-Demolition Measurement (QNDM) Solinas et al. (2023), to efficiently estimate the gradients or the Hessians of a quantum observable. This is a key step and a resource-demanding task when we want to minimize the cost function associated with a quantum observable. The statistical study of the error leads to a further performance improvement with respect to the original proposal in Ref. Solinas et al. (2023). In our detailed analysis, we account for all the resources needed to implement the QNDM approach with a fixed accuracy and compare them to the current state-of-the-art method Mari et al. (2021); Schuld et al. (2019); Cerezo et al. (2021). We find that the QNDM approach is more efficient, i.e. it needs fewer resources, in evaluating the derivatives of a cost function. These advantages are already clear in small dimensional systems and are likely to increase for practical implementations and more realistic situations. A significant outcome of our study is the implementation of the QNDM method in Python, provided in the supplementary material S.Caletti and G.Minuto (2024). Given that most Variational Quantum Algorithms can be formulated within this framework, our results can have significant implications in quantum optimization algorithms and make the QNDM approach a valuable alternative to implement Variational Quantum Algorithms on near-term quantum computers.
@@ I Introduction
We find that, already for cases with limited complexity, the QNDM approach has a substantial advantage over the DM ones. The resource-saving will increase for more practical and interesting problems of intermediate dimension.
@@ II.1 General Settings
This accounts for a multitude of relevant cases such as when $H_{j}$ is a tensor product of any multi-qubits Pauli matrices Mari et al. (2021). However, to further simplify the discussion and the simulation, in the following, we restrict our attention to the case (presented in Fig. 1) in which
@@
To measure the derivative of $f({\vec{\theta}})$ along the $j$-th direction in the parameter space $\vec{\theta}$, we measure $f(\vec{\theta}+s\hat{e}_{j})$ and $f(\vec{\theta}-s\hat{e}_{j})$ where $\hat{e}_{j}$ is the unit vector along the $\theta_{j}$ direction and $s$ is a shift parameter. Then, we calculate the quantity Mari et al. (2021)
$g_{j}=\frac{\partial f({\vec{\theta}})}{\partial\theta_{j}}=\frac{f(\vec{\theta}+s\hat{e}_{j})-f(\vec{\theta}-s\hat{e}_{j})}{2\sin s}\,.$ (7)
With $s=\pi/2$, we obtain the parameter shift rule described in Refs. Li et al. (2017); Mitarai et al. (2018); Schuld et al. (2019); Mitarai and Fujii (2019). Notice that $g_{j}$ is exactly the derivative of $f({\vec{\theta}})$ with no approximation and repeating the procedure for all the directions $\theta_{j}$ we obtain the gradient of $f({\vec{\theta}})$, which is then fed into the classical optimization algorithm.
@@ II.2 Direct Measurement
The standard approach to measure the value of the derivatives of the cost function (7) is the Direct Measurement (DM) method Schuld et al. (2019). This method consists of measuring separately the average values of the operator $\hat{M}$ in the two points, i.e., $\vec{\theta}\pm s\hat{e}_{j}$, and then calculating the derivative as in Eq. (7). To calculate the values of $f(\vec{\theta}+s\hat{e}_{j})$, we run the quantum circuit to implement the corresponding $U(\vec{\theta}+s\hat{e}_{j})$ transformation, and then we perform a projective measure for each Pauli string $\hat{P}_{i}$ that appears in the definition (6) of the observable $\hat{M}$ McArdle et al. (2020); Mari et al. (2021). For every Pauli string $\hat{P}_{i}$ in $\hat{M}$, we have to iterate these steps until the statistical accuracy is reached. The same procedure is then implemented for $f({\vec{\theta}-s\hat{e}_{j}})$ and then, the derivative can be calculated as in Eq. (7).
Following this procedure, a single derivative $g_{j}$ in the DM approach can be written as
$g_{j}=\sum_{i}\frac{h_{i}}{2\sin s}\Big{(}{\rm Tr}_{S}\big{[}\hat{P}_{i}U^{\dagger}(\vec{\theta}+s\hat{e}_{j})\rho_{s}^{0}U(\vec{\theta}+s\hat{e}_{j})\big{]}-{\rm Tr}_{S}\big{[}\hat{P}_{i}U^{\dagger}(\vec{\theta}-s\hat{e}_{j})\rho_{s}^{0}U(\vec{\theta}-s\hat{e}_{j})\big{]}\Big{)}\,,$ (8)
@@
Figure 2: Pictorial representation of the quantum circuit for the implementation of the DM protocol. The $U(\vec{\theta}\pm s\hat{e}_{j})$ transformations are the ones in Eq. (1). The sums represent the fact that the circuits followed by the measure must be run for different shifts ±$\pm s$ and for every Pauli string ($i=1,...,J$) in the observable $\hat{M}$ as in Eq. (6).
@@ II.3 Quantum Non-Demolition Measurement
In this section, we briefly recall the main characteristics of the QNDM method Solinas et al. (2023). The key idea is to store the information about the derivative of the cost function in the phase of an ancillary qubit named the detector. More specifically, by coupling the original system and the detector twice, it is possible to store the information about both $f(\vec{\theta}+s\hat{e}_{j})$ and $f(\vec{\theta}-s\hat{e}_{j})$ in the phase of the detector. Then, the detector is measured to extract information about the derivative Solinas et al. (2023).
@@
Suppose we are interested in the quantum observable $\hat{M}$ in Eq. (6). We add an ancillary qubit and implement the system-detector coupling through the operator $U_{\pm}=\exp\{\pm i\lambda Z_{a}\otimes\hat{M}\}$ where $Z_{a}$ is a Pauli operator acting on the ancillary detector and $\lambda$ is the system-detector coupling constant. We notice that QNDM clearly offers an elegant way to implement the operator $U_{\pm}$, as the product of single Pauli string operators $\prod_{i}^{J}\exp\{\pm i\lambda Z_{a}\otimes h_{i}\hat{P_{i}}\}$ Solinas et al. (2023). Considering the described quantum circuit, this corresponds to the following transformation (in the full system and detector Hilbert space)
$U_{tot}=e^{i\lambda Z_{a}\otimes\hat{M}}U(\vec{\theta}+s\hat{e}_{j})U^{\dagger}(\vec{\theta}-s\hat{e}_{j})e^{-i\lambda Z_{a}\otimes\hat{M}}U(\vec{\theta}-s\hat{e}_{j})\,.$ (9)
We consider as initial state for the system plus the detector $\mathinner{|{\psi_{0}}\rangle}(\mathinner{|{0}\rangle}_{D}+\mathinner{|{1}\rangle}_{D})/\sqrt{2}$. After implementing the transformation (9), we measure the accumulated phase $\exp\{i\phi(\lambda)\}$ on the detector degrees of freedom. It can be shown Solinas et al. (2023); Solinas and Gasparinetti (2015, 2016); Solinas et al. (2021, 2022) that this is a quasi-characteristic function $\mathcal{G}_{\lambda}$
@@
where $\mathinner{|{0}\rangle}_{D}$ and $\mathinner{|{1}\rangle}_{D}$ are eigenstates of the detector operator $Z_{a}$.
Figure 3: Quantum circuit of the implementation of the QNDM protocol. Here, $H$ is the Hadamard gate, $S$ is the phase gate, $U_{1}=U({\vec{\theta}}-s\hat{e}_{j})$, $U_{2}=U^{\dagger}({\vec{\theta}}-s\hat{e}_{j})U({\vec{\theta}}+s\hat{e}_{j})$ and $U_{\pm}=\exp\{\pm i\lambda\hat{Z_{a}}\otimes\hat{M}\}$ is the system-detector coupling operator.
Taking the derivatives of $\mathcal{G}_{\lambda}$ with respect to $\lambda$ and evaluating them in $\lambda=0$, we have access to the information of the different moments of the distribution describing the variation of the cost function Solinas et al. (2023); Solinas and Gasparinetti (2015, 2016); Solinas et al. (2021, 2022). Here, we are interested only in the first moment which directly gives us $g_{j}$. Following Ref. Solinas et al. (2023), it can be shown that the first derivative of $\mathcal{G}_{\lambda}$ reads
$-i\partial_{\lambda}\mathcal{G}_{\lambda}\Big{|}_{\lambda=0}$ $=$ $2{\rm Tr}_{S}[U^{\dagger}(\vec{\theta}+s\hat{e}_{j})\hat{M}U(\vec{\theta}+s\hat{e}_{j})\rho_{S}^{0}-U^{\dagger}(\vec{\theta}-s\hat{e}_{j})\hat{M}U(\vec{\theta}-s\hat{e}_{j})\rho_{S}^{0}]$ (11)
$=$ $2\sum_{i}h_{i}{\rm Tr}_{S}[U^{\dagger}(\vec{\theta}+s\hat{e}_{j})\hat{P_{i}}U(\vec{\theta}+s\hat{e}_{j})\rho_{S}^{0}-U^{\dagger}(\vec{\theta}-s\hat{e}_{j})\hat{P_{i}}U(\vec{\theta}-s\hat{e}_{j})\rho_{S}^{0}]\,,$
@@
$-i\partial_{\lambda}\mathcal{G}_{\lambda}\Big{|}_{\lambda=0}=2\sin(s)g_{j}.$ (12)
@@
In practice, it is convenient to approximate the derivative of the quasi-characteristic function for small $\lambda$, i.e. in the linear regime. In this approximation, the derivative is given by Solinas et al. (2023)
$-i\partial_{\lambda}\mathcal{G}_{\lambda}|_{\lambda=0}=\partial_{\lambda}\phi(\lambda)|_{\lambda=0}\approx\frac{\phi(\lambda)-\phi(0)}{\lambda}=\frac{\phi(\lambda)}{\lambda}$ (13)
@@
$g_{j}\sim-\frac{\arcsin(2P_{0}-1)}{2\lambda\sin s}\,,$ (16)
@@ III Mean Squared Error and cost analysis
Understanding how statistical errors affect the estimate of the derivatives and gradient plays a pivotal role in judging the performance and reliability of the different methods presented above. As a measure of the quality of the estimation of the cost function derivative, we consider Mean Squared Error ($MSE$), i.e., the average of the difference between the estimated values $\hat{g}_{j}$ and the exact value $g_{j}$, squared. We introduce the bias and variance as Cox and Hinkley (1974)
$\text{Bias}(\hat{g}_{j})=\mathbb{E}(\hat{g}_{j})-g_{j}\,,$ (17)
$\text{Var}(\hat{g}_{j})=\mathbb{E}[(\hat{g}_{j}-\mathbb{E}(\hat{g}_{j}))^{2}]\,,$ (18)
where $\mathbb{E}(\cdot)$ is the expectation over the statistical distribution of the measurement outcomes. The bias represents a systematic distortion of a statistical result, and it remains present even in the limit of an infinite number of shots $N$. On the other hand, variance quantifies the degree of dispersion, indicating how far a collection of values deviates from its mean value. Unlike the bias, the variance depends on the shot number $N$ Mari et al. (2021). In terms of these quantities, we define the mean square error (MSE) as
$\text{MSE}(\hat{g}_{j})=\text{Bias}(\hat{g}_{j})^{2}+\text{Var}(\hat{g}_{j})\,.$ (19)
@@
$\text{Bias}_{QNDM}(\hat{g}_{j})=-i\left(\frac{\phi(\lambda)-\phi(0)}{\lambda}-\partial_{\lambda}\mathcal{G}_{\lambda}\right)=\frac{\lambda\partial^{2}_{\lambda}\mathcal{G}_{\lambda}}{2}+\mathcal{O}(\lambda^{2})\,,$ (20)
$\text{Var}_{QNDM}(\hat{g}_{j})=\sigma_{QNDM}^{2}\,.$ (21)
@@
$\text{MSE}_{QNDM}(\hat{g}_{j})=\frac{\lambda^{2}(\partial^{2}_{\lambda}\mathcal{G}_{\lambda})^{2}}{4}+\frac{\sigma^{2}_{D}}{4N_{QNDM}\sin^{2}s}\frac{1}{\lambda^{2}(1-(2P_{0}-1)^{2})}.$ (23)
Notably, in MSEQNDM, the variance term is directly proportional to $\lambda^{-2}$ while the bias term is proportional to $\lambda^{2}$. Thus, the optimal reduction of the MSE hinges on selecting an appropriate value for $\lambda$. We determined that an optimal balance is achieved for $\lambda=0.1$. With this specific choice of $\lambda$, in the shot regime commonly employed in optimization, $N\in[10^{3}-10^{4}]$ Kandala et al. (2017); Schuld and Killoran (2019), $\sigma^{2}_{QNDM}$ dominates over the bias squared term so that $\text{MSE}_{QNDM}(\hat{g}_{j})\approx\text{Var}_{QNDM}(\hat{g}_{j})=\sigma^{2}_{QNDM}$. Consequently, in the computational section IV, we exclusively consider the contribution from the variance to calculate the MSEQNDM error.
In the DM approach, we need to measure the system qubits twice to obtain $f(\vec{\theta}+s\hat{e}_{j})$ and $f(\vec{\theta}-s\hat{e}_{j})$. Proceeding as above, we have
$\text{Bias}_{DM}(\hat{g}_{j})=0$ (24)
$\text{Var}_{DM}(\hat{g}_{j})=\frac{\sigma_{DM}^{2}(\vec{\theta}+s\hat{e}_{j})+\sigma_{DM}^{2}(\vec{\theta}-s\hat{e}_{j})}{4\sin^{2}s}.$ (25)
The bias contribution for DM vanishes because the parameter shift rule calculates the exact value of derivative Mari et al. (2021). The variance, instead, is composed by the terms $\sigma_{DM}^{2}(\vec{\theta}\pm s\hat{e}_{j})$. Intuitively, they correspond to the evaluation of $f(\vec{\theta}\pm s\hat{e}_{j})$. Following Ref. Mari et al. (2021), we assume that the variance of the measured observable depends weakly on the parameter shift such that $\sigma_{DM}^{2}(\vec{\theta}+s\hat{e}_{j})+\sigma_{DM}^{2}(\vec{\theta}-s\hat{e}_{j})\sim 2\sigma_{DM}^{2}$ for any value of $s$. As we have seen in sec II.2, to obtain the derivatives, we have to measure each Pauli string individually in the parameter space point $\vec{\theta}\pm s\hat{e}_{j}$ (see Eq. (8) and Mari et al. (2021); Cerezo et al. (2021)). Then, the statistical error formula for the estimation of $f(\vec{\theta})$ is given by
@@
$\text{MSE}_{DM}(\hat{g}_{j})=\sum_{i=1}^{J}\frac{h^{2}_{i}\sigma_{s}^{2}}{2N_{DM}\sin^{2}s}.$ (27)
An important remark must be made about the results obtained in Eq. (23) and (27). The mean square error MSEQNDM does not depend on the number of the Pauli string in $\hat{M}$ or their weights, while MSEDM does. The reason for this is tracked back to the different ways to extract information about the cost function $f$ and its derivatives. In the QNDM protocol all the information about $\hat{M}$ (and the Pauli strings) are stored in series through the system-detector coupling. On the contrary, in the DM protocol, the information about the Pauli strings is retrieved one at a time by running the circuits and performing the measurements. We can estimate exactly the number of shots to get the same precision in MSEDM and MSEQNDM by using Eqs. (23) and (27). We obtain
$\frac{\text{MSE}_{DM}(\hat{g}_{j})}{\text{MSE}_{QNDM}(\hat{g}_{j})}=1\longrightarrow N_{DM}=N_{QNDM}\frac{2\sigma_{s}^{2}\lambda^{2}(1-(2P_{0}-1)^{2})}{\sigma^{2}_{D}}\sum_{i}h^{2}_{i},$ (28)
where the $\sum_{i}h^{2}_{i}$ grows proportionally to the number of Pauli strings.

### v3

@@ Abstract
We present a detailed numerical study of an alternative approach, named Quantum Non-Demolition Measurement (QNDM) Solinas et al. (2023), to efficiently estimate the gradients or the Hessians of a quantum observable. This is a key step and a resource-demanding task when we want to minimize the cost function associated with a quantum observable. In our detailed analysis, we account for all the resources needed to implement the QNDM approach with a fixed accuracy and compare them to the current state-of-the-art method Mari et al. (2021); Schuld et al. (2019); Cerezo et al. (2021). We find that the QNDM approach is more efficient, i.e. it needs fewer resources, in evaluating the derivatives of a cost function. These advantages are already clear in small dimensional systems and are likely to increase for practical implementations and more realistic situations. A significant outcome of our study is the implementation of the QNDM method in Python, provided in the supplementary material S.Caletti and G.Minuto (2024). Given that most Variational Quantum Algorithms can be formulated within this framework, our results can have significant implications in quantum optimization algorithms and make the QNDM approach a valuable alternative to implement Variational Quantum Algorithms on near-term quantum computers.
@@ I Introduction
We find that, already for cases with limited complexity, the QNDM approach has a substantial advantage over the DM ones. The resource-saving will increase for more practical and interesting problems of intermediate dimension. We observe that there are other methods that rely on measuring an additional detector, such as the approach proposed in Guerreschi and Smelyanskiy (2017). However, the computational cost of obtaining the derivative with this method is comparable to that of the DM approach.
@@ II.1 General Settings
This accounts for a multitude of relevant cases such as when $\hat{H}_{j}$ is a tensor product of any multi-qubits Pauli matrices Mari et al. (2021). However, to further simplify the discussion and the simulation, in the following, we restrict our attention to the case (presented in Fig. 1) in which
@@
To measure the derivative of $f({\vec{\theta}})$ along the $l$-th direction in the parameter space $\vec{\theta}$, we measure $f(\vec{\theta}+s\hat{e}_{l})$ and $f(\vec{\theta}-s\hat{e}_{l})$ where $\hat{e}_{l}$ is the unit vector along the $\theta_{l}$ direction and $s$ is a shift parameter. Then, we calculate the quantity Mari et al. (2021)
$g_{l}=\frac{\partial f({\vec{\theta}})}{\partial\theta_{l}}=\frac{f(\vec{\theta}+s\hat{e}_{l})-f(\vec{\theta}-s\hat{e}_{l})}{2\sin s}\,.$ (7)
Notice that $g_{l}$, unlike finite difference methods, is not an approximation but the exact derivative of $f({\vec{\theta}})$. By repeating this procedure for all directions $\theta_{l}$ we obtain the full gradient of $f({\vec{\theta}})$, which is then fed into the classical optimization algorithm. Setting $s=\pi/2$ yields the parameter shift rule described in Refs. Li et al. (2017); Mitarai et al. (2018); Schuld et al. (2019); Mitarai and Fujii (2019).
@@ II.2 Direct Measurement
The standard approach to measure the value of the derivatives of the cost function (7) is the Direct Measurement (DM) method Schuld et al. (2019). This method consists of measuring separately the average values of the operator $\hat{M}$ in the two points, i.e., $\vec{\theta}\pm s\hat{e}_{l}$, and then calculating the derivative as in Eq. (7). To calculate the values of $f(\vec{\theta}+s\hat{e}_{l})$, we run the quantum circuit to implement the corresponding $U(\vec{\theta}+s\hat{e}_{l})$ transformation, and then we perform a projective measurement for each Pauli string $\hat{P}_{i}$ that appears in the definition (6) of the observable $\hat{M}$ McArdle et al. (2020); Mari et al. (2021). For every Pauli string $\hat{P}_{i}$ in $\hat{M}$, we have to iterate these steps until the statistical accuracy is reached. The same procedure is then implemented for $f({\vec{\theta}-s\hat{e}_{l}})$ and then, the derivative can be calculated as in Eq. (7).
Following this procedure, a single derivative $g_{l}$ in the DM approach can be written as
$g_{l}=\sum_{i}\frac{h_{i}}{2\sin s}\Big{(}{\rm Tr}_{S}\big{[}\hat{P}_{i}U^{\dagger}(\vec{\theta}+s\hat{e}_{l})\rho_{s}^{0}U(\vec{\theta}+s\hat{e}_{l})\big{]}-{\rm Tr}_{S}\big{[}\hat{P}_{i}U^{\dagger}(\vec{\theta}-s\hat{e}_{l})\rho_{s}^{0}U(\vec{\theta}-s\hat{e}_{l})\big{]}\Big{)}\,,$ (8)
@@
Figure 2: Pictorial representation for calculation of derivative of the function $f({\vec{\theta}})={\mathinner{\langle{0}|{U^{\dagger}({\vec{\theta}})\hat{M}U({\vec{\theta}})}|{0}\rangle}}$ with $\hat{M}=\sum_{i=1}^{J}h_{i}\hat{P}_{i}$ using the DM protocol, as described in Eq. 8. The $n$ qubits circuit is is evolved under the parametric transformation $U(\vec{\theta}+s\hat{e}_{l})$, defined similarly to Eq.1, followed by a base rotation $P_{i}$, which represents the $i$-th tensor product of Pauli matrices. A global measurement is then performed on the register, yielding a real-valued output. The first summation indicates that the circuit must be executed for each Pauli string ($i=1,\ldots,J$) in the observable $\hat{M}$. The second summation accounts for the execution of the circuit with different parameter shifts ±$\pm s$, as shown in Eq. (8).
@@ II.3 Quantum Non-Demolition Measurement
In this section, we briefly recall the main characteristics of the QNDM method Solinas et al. (2023). The key idea is to store the information about the derivative of the cost function in the phase of an ancillary qubit named the detector. More specifically, by coupling the original system and the detector twice, it is possible to store the information about both $f(\vec{\theta}+s\hat{e}_{l})$ and $f(\vec{\theta}-s\hat{e}_{l})$ in the phase of the detector. Then, the detector is measured to extract information about the derivative Solinas et al. (2023).
@@
Suppose we are interested in the quantum observable $\hat{M}$ in Eq. (6). We add an ancillary qubit and implement the system-detector coupling through the operator $U_{\pm}=\exp\{\pm i\lambda Z_{a}\otimes\hat{M}\}$ where $Z_{a}$ is a Pauli operator acting on the ancillary detector and $\lambda$ is the system-detector coupling constant. We notice that if $\hat{M}$ is a complex operator, the implementation of operator $U_{\pm}$ is, in general, very resource-consuming. This might seem a critical drawback, but the QNDM approach offers a clear and elegant way to bypass this problem, instead of implementing the operator $U_{\pm}=\exp\{\pm i\lambda Z_{a}\otimes\sum_{i=1}^{J}h_{i}\hat{P}_{i}\}$, we can implement the product of single Pauli string operators $\prod_{i}^{J}\exp\{\pm i\lambda Z_{a}\otimes h_{i}\hat{P_{i}}\}$ without relying on the Trotterization approximation Hatano and Suzuki (2005), as demonstrated in Solinas et al. (2023). Considering the described quantum circuit, this corresponds to the following transformation (in the full system and detector Hilbert space)
$U_{tot}=e^{i\lambda Z_{a}\otimes\hat{M}}U(\vec{\theta}+s\hat{e}_{l})U^{\dagger}(\vec{\theta}-s\hat{e}_{l})e^{-i\lambda Z_{a}\otimes\hat{M}}U(\vec{\theta}-s\hat{e}_{l})\,.$ (9)
The initial state for the system plus the detector $\mathinner{|{\psi_{0}}\rangle}(\mathinner{|{0}\rangle}_{D}+\mathinner{|{1}\rangle}_{D})/\sqrt{2}$, where a Hadamard gate is applied to the initial state of the detector. After implementing the transformation (9), we measure the accumulated phase $\exp\{i\phi(\lambda)\}$ on the detector degrees of freedom. It can be shown Solinas et al. (2023); Solinas and Gasparinetti (2015, 2016); Solinas et al. (2021, 2022) that this is a quasi-characteristic function $\mathcal{G}_{\lambda}$
@@
where $\mathinner{|{0}\rangle}_{D}$ and $\mathinner{|{1}\rangle}_{D}$ are eigenstates of the detector operator $Z_{a}$, while $\rho_{D}^{0}$ and $\rho_{D}^{f}$ are the density matrices of the detector before and after the application of $U_{tot}$, respectively.
Figure 3: Pictorial representation of the quantum circuit implementing the QNDM protocol. Here, $H$ is the Hadamard gate, $S$ is the phase gate, $U_{1}=U({\vec{\theta}}-s\hat{e}_{l})$, $U_{2}=U^{\dagger}({\vec{\theta}}-s\hat{e}_{l})U({\vec{\theta}}+s\hat{e}_{l})$ and $U_{\pm}=\exp\{\pm i\lambda\hat{Z_{a}}\otimes\hat{M}\}$ is the system-detector coupling operator.
Taking the derivatives of $\mathcal{G}_{\lambda}$ with respect to $\lambda$ and evaluating them in $\lambda=0$, we have access to the information of the different moments of the distribution describing the variation of the cost function Solinas et al. (2023); Solinas and Gasparinetti (2015, 2016); Solinas et al. (2021, 2022). Here, we are interested only in the first moment which directly gives us $g_{l}$. Following Ref. Solinas et al. (2023), it can be shown that the first derivative of $\mathcal{G}_{\lambda}$ reads
$-i\partial_{\lambda}\mathcal{G}_{\lambda}\Big{|}_{\lambda=0}$ $=$ $2{\rm Tr}_{S}[U^{\dagger}(\vec{\theta}+s\hat{e}_{l})\hat{M}U(\vec{\theta}+s\hat{e}_{l})\rho_{S}^{0}-U^{\dagger}(\vec{\theta}-s\hat{e}_{l})\hat{M}U(\vec{\theta}-s\hat{e}_{l})\rho_{S}^{0}]$ (11)
$=$ $2\sum_{i}h_{i}{\rm Tr}_{S}[U^{\dagger}(\vec{\theta}+s\hat{e}_{l})\hat{P_{i}}U(\vec{\theta}+s\hat{e}_{l})\rho_{S}^{0}-U^{\dagger}(\vec{\theta}-s\hat{e}_{l})\hat{P_{i}}U(\vec{\theta}-s\hat{e}_{l})\rho_{S}^{0}]\,,$
@@
$-i\partial_{\lambda}\mathcal{G}_{\lambda}\Big{|}_{\lambda=0}=2\sin(s)g_{l}.$ (12)
@@
We note that it is possible to approximate, at the linear regime, the derivative of the quasi-characteristic function to the argument of the accumulated detector phase $\phi(\lambda)$Solinas et al. (2023). In practice, this approximation holds if the coefficient in the argument of the system-detector coupling operator, $U_{\pm}=\exp\{\pm i\lambda\hat{Z_{a}}\otimes\hat{M}\}$, is small. This condition is satisfied for sufficiently small $\lambda$, allowing the derivative to be expressed as
$-i\partial_{\lambda}\mathcal{G}_{\lambda}|_{\lambda=0}=-i\partial_{\lambda}e^{i\phi(\lambda)}|_{\lambda=0}=\partial_{\lambda}\phi(\lambda)|_{\lambda=0}\approx\frac{\phi(\lambda)-\phi(0)}{\lambda}=\frac{\phi(\lambda)}{\lambda}$ (13)
@@
$g_{l}\sim-\frac{\arcsin(2P_{0}-1)}{2\lambda\sin s}\,,$ (16)
@@ III Mean Squared Error and cost analysis
Understanding how statistical errors affect the estimate of the derivatives and gradient plays a pivotal role in judging the performance and reliability of the different methods presented above. As a measure of the quality of the estimation of the cost function derivative, we consider Mean Squared Error ($MSE$), i.e., the average of the difference between the estimated values $\hat{g}_{l}$ and the exact value $g_{l}$, squared. We introduce the bias and variance as Cox and Hinkley (1974)
$\text{Bias}(\hat{g}_{l})=\mathbb{E}(\hat{g}_{l})-g_{l}\,,$ (17)
$\text{Var}(\hat{g}_{l})=\mathbb{E}[(\hat{g}_{l}-\mathbb{E}(\hat{g}_{l}))^{2}]\,,$ (18)
where $\mathbb{E}(\cdot)$ is the expectation over the statistical distribution of the measurement outcomes. The bias represents a systematic distortion of a statistical result, and it remains present even in the limit of an infinite number of shots $N$. This is a consequence of the linear approximation used in Eq. (13). On the other hand, variance quantifies the degree of dispersion, indicating how far a collection of values deviates from its mean value. Unlike the bias, the variance depends on the shot number $N$ Mari et al. (2021). In terms of these quantities, we define the mean square error (MSE) as
$\text{MSE}(\hat{g}_{l})=\text{Bias}(\hat{g}_{l})^{2}+\text{Var}(\hat{g}_{l})\,.$ (19)
@@
$\text{Bias}_{QNDM}(\hat{g}_{l})=\left(\frac{\phi(\lambda)}{\lambda}-i\partial_{\lambda}\mathcal{G}_{\lambda}\right)=\frac{\lambda\partial^{2}_{\lambda}\mathcal{G}_{\lambda}}{2}+\mathcal{O}(\lambda^{2})\,,$ (20)
$\text{Var}_{QNDM}(\hat{g}_{l})=\sigma_{QNDM}^{2}\,.$ (21)
@@
$\text{MSE}_{QNDM}(\hat{g}_{l})=\frac{\lambda^{2}(\partial^{2}_{\lambda}\mathcal{G}_{\lambda})^{2}}{4}+\frac{\sigma^{2}_{D}}{4N_{QNDM}\sin^{2}s}\frac{1}{\lambda^{2}(1-(2P_{0}-1)^{2})}.$ (23)
Notably, in MSEQNDM, the variance term is directly proportional to $\lambda^{-2}$ while the bias term is proportional to $\lambda^{2}$. Thus, the optimal reduction of the MSE hinges on selecting an appropriate value for $\lambda$.
As highlighted in the previous section, the derivative of the quasi-characteristic can be approximately calculated if the coefficient in the argument of the system-detector coupling operator, $U_{\pm}=\exp\{\pm i\lambda\hat{Z_{a}}\otimes\hat{M}\}$, is small. Consequently, the appropriate choice of $\lambda$ depends, also, from the coefficients $h_{i}$ of observable $\hat{M}=\sum_{i=1}^{J}h_{i}\hat{P}_{i}$. In the computational analysis, section IV, we present a significant example where a suitable value of $\lambda$ is selected.
In the DM approach, we need to measure the system qubits twice to obtain $f(\vec{\theta}+s\hat{e}_{l})$ and $f(\vec{\theta}-s\hat{e}_{l})$. Proceeding as above, we have
$\text{Bias}_{DM}(\hat{g}_{l})=0$ (24)
$\text{Var}_{DM}(\hat{g}_{l})=\frac{\sigma_{DM}^{2}(\vec{\theta}+s\hat{e}_{l})+\sigma_{DM}^{2}(\vec{\theta}-s\hat{e}_{l})}{4\sin^{2}s}.$ (25)
The bias contribution for DM vanishes because the parameter shift rule calculates the exact value of derivative Mari et al. (2021). The variance, instead, is composed by the terms $\sigma_{DM}^{2}(\vec{\theta}\pm s\hat{e}_{l})$. Intuitively, they correspond to the evaluation of $f(\vec{\theta}\pm s\hat{e}_{l})$. Following Ref. Mari et al. (2021), we assume that the variance of the measured observable depends weakly on the parameter shift such that $\sigma_{DM}^{2}(\vec{\theta}+s\hat{e}_{l})+\sigma_{DM}^{2}(\vec{\theta}-s\hat{e}_{l})\sim 2\sigma_{DM}^{2}$ for any value of $s$. As we have seen in sec II.2, to obtain the derivatives, we have to measure each Pauli string individually in the parameter space point $\vec{\theta}\pm s\hat{e}_{l}$ (see Eq. (8) and Mari et al. (2021); Cerezo et al. (2021)). Then, the statistical error formula for the estimation of $f(\vec{\theta})$ is given by
@@
$\text{MSE}_{DM}(\hat{g}_{l})=\sum_{i=1}^{J}\frac{h^{2}_{i}\sigma_{s}^{2}}{2N_{DM}\sin^{2}s}.$ (27)
A remark must be made regarding the results obtained in Eqs. (23) and (27). As demonstrated in section IV, an appropriate choice of $\lambda$ could provide a constant advantage in terms of error, giving to QNDM an advantage in terms of efficiency cost.
$\frac{\text{MSE}_{DM}(\hat{g}_{l})}{\text{MSE}_{QNDM}(\hat{g}_{l})}>1$ (28)